OpenStudy (anonymous):

For the line that passes through the points (0, 2) and (-6, 0), the x-intercept is_______________and the y-intercept is _______________________________

OpenStudy (anonymous):

I think it is, x= -6, and Y= 2... Is that right?

pooja195 (pooja195):

We need to figure out the slope first do you know how to do that?

OpenStudy (anonymous):

To be honest, I don't think we need to go through that.... I mean I just wanted someone to check my work, because if I got it wrong, I need someone to explain it to me then....

pooja195 (pooja195):

Thats wrong.

OpenStudy (anonymous):

Okay, I do know how to find the slope.

OpenStudy (anonymous):

I just don't have graph paper.

pooja195 (pooja195):

You dont need graph paper.

pooja195 (pooja195):

Use the slope Formula \[\huge~\rm~\frac{ y_2-y_1 }{ x_2-x_1}\]

OpenStudy (anonymous):

I never used that formula before? Are you sure it is right?

pooja195 (pooja195):

Its the slope formula.

OpenStudy (anonymous):

I used y=mx+b for the slope, and on the graph paper, I used to draw the line on the squares, to get y=mx+b... i'm in 8th grade..... so... idk... but actually it is right... I turned it in to my teacher and got it right.

OpenStudy (anonymous):

x-intercept = -6 and y-intercept= 2

pooja195 (pooja195):

Thats weird. Because thats the slope formula and to find the link you plug it into the Point slope equations and then solve and get a y=mx+b answer but whatever :)

OpenStudy (anonymous):

Okay.